Homework 2 11-721: Grammars and Lexicons, Fall 2008 Due Monday, October 6, 2008

1 Based on Radford, Exercise XVI, page 225

Assume the following phrase structure rules:

S --> N''  VP
VP --> V'
VP --> V' PP
V' --> V
V' --> V N''
PP --> P  N''
N'' --> (Det) N'
N' --> N' PP
N' --> N  (PP)
N --> N CONJ N
N' --> N' CONJ N'
N'' --> N'' CONJ N''
Det --> that
Det --> his
Det --> the
CONJ --> and
V --> like
VP --> just don't get on
N --> picture
N --> pushchair
N --> workers
N --> managers
N --> factory
N'' --> I
N'' --> him
P --> of
P --> in

Question 1: Note that I and him are N'' and not N. Write an ungrammatical sentence that would be generated if “I” and “him” were N and not NP. This grammar generates a lot of ungrammatical sentences, so be sure to write one that is ungrammatical for the right reason. Question 2: Note that his is a determiner. Write an ungrammatical sentence that would be generated if his were an N. Write a sentence that could not be generated if his were an NP. Question 3: Assuming that there is an attachment ambiguity in that picture of him in his pushchair, what are the two meanings? Question 4: The phrase structure rules don’t generate a tree that corresponds to one of the meanings. Which one is it? Question 5: Draw a phrase structure tree for the meaning that is generated. Question 6: There is also an attachment ambiguity in The workers and the managers in the factory just don’t get on. What are the two meanings? Question 7: The phrase structure rules don’t generate a tree that corresponds to one of the meanings. Which one is it? Question 8: Draw a phrase structure tree for the meaning that is generated. Question 9: Radford suggests to reformulate the rule that introduces PP adjuncts into three rules, one for adjuncts to X, one for adjuncts to X’, and one for adjuncts to X”. Write the new phrase structure rules. Question 10: Using the new rules, draw a tree for the meaning of that picture of him in his pushchair that wasn’t previously generated. Question 11: Using the new rules, draw a tree for the meaning of The workers and the managers in the factory just don’t get on that wasn’t previously generated.

2 Possessive Noun Phrases

In the last problem, possessive pronouns like his were treated as determiners. However, possessors can be full noun phrases as in the boy across the street’s toy where the possessor is the boy across the street.

The ’s is a clitic, not a full word, but not a prefix or suffix. One of the things that makes it different from prefixes or suffixes is that it attaches to words that aren’t related to it. In the boy across the street’s toy the ’s indicates that the boy is a possessor, but it attaches to street. In the phrase structure rules below, the ’s is treated as a terminal symbol that is a daughter of NP.

Task 1: Using the phrase structure rules below, draw two trees for the wife of your friend’s brother, which has two different meanings. Assume that of your friend is an argument, not an adjunct. For the sake of simplicity, your is still treated as a determiner because it does not occur with ’s.

Task 2: Using the phrase structure rules below, draw two trees for We gave her dog biscuits. The verb give should have three arguments: agent, theme, and recipient.

S --> N''  VP
VP -->  V  N''  N''
PP -->  P  N''
N'' --> (Det) N'
N'  -->  N  (PP)
N'' -->  NP  's   N'
In the previous rule, the N'' has three daughters, NP, 's, and N'.
Det --> the
Det --> your
Det --> her
N'' --> her
N'' --> we
N -->  N   N

Task 3 (Extra Credit, five percent): When there is an NP consisting of NP 's N', what principle determines that the N-bar and not the possessor NP is the head of the whole NP? I’m looking for a syntactic principle, which might be found somewhere in Radford’s book. A semantic principle is reasonable — the boy’s book is a book, not a boy — but that is not the answer I’m looking for. The answer is very specific and pretty short.

3 Drawing F-structures

On the next page there is a lexicon and a reduced set of rules in LFG notation. Draw the f-structures for the two readings of the friend of your wife’s brother.

PP →  P  N′′
↑=↓  ↑=↓

N′′ →  (Det)  N′
↑=↓  ↑=↓

N′′ →  (Det)  N′
(↑ poss)= ↓  ↑=↓
(↓ pron)=c +

N′ →  N  (PP)
↑=↓  (↑ obl-poss) = ↓

N′′ →  N′′  ′s  N′
(↑ POSS) = ↓  ↑=↓

the  det (↑ def) = +
wife  n  (↑ pred) = ′wife′
of  P
your  det (↑ pred) = ′pro′
(↑ num) = sg
(↑ pers) = 3
(↑ pron) = +
friend n  (↑ pred) = ′friend′
brother n  (↑ pred) = ′brother′